OpenStudy (anonymous):

Need math help!! (Question in comments because of how the question is formed)

OpenStudy (anonymous):

Which quadratic equation has the solutions? \[x=\frac{ 6\pm \sqrt{22} }{ 7}\] A. -4x^2+5x+2=0 B. 5x^2-13x+7=0 C. x^2+7x+10=0 D. 7x^2-12x+2=0

OpenStudy (anonymous):

@Nnesha Can you help ???